### v2

# MAROON: A Framework for the Joint Characterization of Near-Field High-Resolution Radar and Optical Depth Imaging Techniques

Abstract.
Utilizing the complementary strengths of wavelength-specific range or depth sensors is crucial for robust computer-assisted tasks such as autonomous driving. Despite this, there is still little research done at the intersection of optical depth sensors and radars operating close range, where the target is decimeters away from the sensors. Together with a growing interest in high-resolution imaging radars operating in the near field, the question arises how these sensors behave in comparison to their traditional optical counterparts. In this work, we take on the unique challenge of jointly characterizing depth imagers from both, the optical and radio-frequency domain using a multimodal spatial calibration. We collect data from four depth imagers, with three optical sensors of varying operation principle and an imaging radar. We provide a comprehensive evaluation of their depth measurements with respect to distinct object materials, geometries, and object-to-sensor distances. Specifically, we reveal scattering effects of partially transmissive materials and investigate the response of radio-frequency signals. All object measurements will be made public in form of a multimodal dataset, called MAROON.
Time of Flight, Radar Imaging, Radio Frequency, mmWave Imaging, MIMO Radar, Depth Camera, Spatial Calibration, Multimodal Sensor Fusion, Dataset
(Sec. 5.1)
(Sec. 5.2)
(Sec. 6)
Figure 1. Recent developments for near-field imaging radars enabled the acquisition of high-resolution depth images, and the sensors are now increasingly gaining attention as complementary modalities to optical depth sensing. Direct comparisons from our MAROON dataset, however, highlight significant differences between radar and optical reconstructions. This work employs the collected multimodal data of four depth imagers, depicted on the left, to systematically characterize these fundamental differences together with sensor-specific findings in a joint evaluation framework.
1. Introduction
Real-world computer-assisted tasks, for instance in robotics and tracking applications, frequently require the immediate assessment of spatial information to accurately reason about the environment at a specific point in time, which has led to the development of several single-view range and depth sensors. For autonomous driving, it has been shown that utilizing multimodal depth sensing techniques from both, the optical (lidar) and radio-frequency (radar) domain, can lead to superior performance and robustness in computer-assisted tasks (Velasco-Hernandez et al., 2020). Due to its environment, the autonomous driving industry has traditionally concentrated on far-field range sensing, with an unambiguous range of several meters and beyond. As recent high-resolution radio-frequency technologies utilize the concept of radar imaging to produce 3D information in form of a depth map — similar to optical depth or RGB-D cameras — they also become more popular in close range, where the target of interest is up to a few decimeters away from the sensor; however, a comprehensive and detailed characterization of these radar imaging technologies, which frequently operate in the radar’s near field, has yet to be realized.
As part of this work, we devised a dataset MAROON (Multimodal Aligned Radio and Optical frequency Object Reconstructions in the Near Field) (cf. Section 5) that enables studying of different sensor modalities in direct comparison. As is immediately visible in Figure 1 (left), the reconstructions of near-field imaging radars appear fundamentally different in comparison to their well-researched counterparts in the optical domain.
A key advantage of radar is that it is insensitive to environmental light and can penetrate, for instance, fabric and dust. Following the success of Google’s project Soli (Lien et al., 2016) for gesture sensing, radars were utilized in close range for the detection of vital signs (Vilesov et al., 2022), activity recognition (Braeunig et al., 2023), people tracking (Zewge et al., 2019) and human body reconstruction (Chen et al., 2022; Chen et al., 2023). With the growing trend towards larger antenna apertures to achieve high-resolution imaging (Schwarz et al., 2022; Chen et al., 2023), radars will more frequently operate in the near field, as determined by the Fraunhofer boundary condition (Selvan and Janaswamy, 2017). At the same time, characteristics of near-field radar are generally under-researched.
Figure 2. Example data of the Plunger object from the MAROON dataset. In the upper left, all reconstructions are spatially aligned with respect to the RF ToF coordinate system. The RF ToF colorscale encodes the normalized reconstruction confidence (cf. Section 4.2.2).
Drawing on prior research about wavelength-specific strengths and weaknesses, this paper addresses the unique challenge of characterizing various optical depth-imaging techniques alongside a high-resolution multiple-input multiple-output (MIMO) imaging radar in the near field. The latter is interchangeably referred to as a radio-frequency (RF) Time-of-Flight (ToF) depth imager. To this end, we mutually calibrated sensors of four different depth sensing technologies, that is active and passive stereo, near-infrared (NIR) ToF, and RF ToF in the millimeter-wave range.
Using these sensors, we captured the comprehensive MAROON dataset of various household objects and construction materials, of which example data is shown in Figure 2. The dataset enables an exploration of sensor-specific characteristics with respect to various object materials, geometries and distances from the sensor. Furthermore, the spatial calibration enables a direct comparison of all sensor reconstructions in a common metrical space. We then compare these results to a reference ground truth reconstruction from a state-of-the-art multi-view stereo (MVS) system. To achieve this, we developed a joint sensor evaluation framework that measures reconstruction differences between sensors and ground truth using different metrics, and that provides supplementary visualizations tailored to identify sensor-specific trends across multiple objects. Our findings reveal fundamental differences between optical and RF sensors, as well as characteristics arising from specific depth measurement techniques within a single frequency domain. We detail sensor-specific characteristics based on several key insights. First, by assessing deviation from the ground truth across different objects and object-to-sensor distances, we identify common sensor-specific trends. Given these, RF ToF reconstructions are found to be less complete compared to optical sensors. Lastly, we demonstrate that ToF sensors exhibit systematic depth errors resulting from interactions within partially transmissive media, such as subsurface scattering.
To summarize, our contributions are the following:
  • A framework enabling the systematic comparative characterization of depth-sensing techniques in both, the optical and radio-frequency signal domain.
  • A novel multimodal dataset, MAROON, comprising common objects in the near field, which were captured with the unique setup of four jointly calibrated depth sensors.
  • We publicly release this dataset together with the raw radar measurements, thus enabling the community to explore different signal reconstruction and filtering techniques. [^1] State-of-the-art methods for reconstruction and filtering will be publicly released together with the dataset viewer upon acceptance.
  • A detailed analysis of trends and sensor-specific effects emerging from that dataset. This includes aspects of different object materials, geometries, and distances to the sensors, signal response and reconstruction quality of imaging radars, as well as ToF scattering effects of partially transmissive materials.
2. Related Work
While a considerable amount of literature exists on optical and RF depth sensors in isolation, no directly related work on the joint characterization of these two domains has been identified. Instead, the first two sections comprise an overview of existing research about sensor characteristics self-contained within a single frequency domain. We further address the sensor fusion of optical and RF sensors, since in that research direction the complementary strengths of the sensors are utilized as well.
2.1. Optical Depth Sensing
Depth cameras have been characterized with respect to a number of different aspects, and related work can be broadly classified into three categories: the sensor technologies, the capture environments, and the methods of comparison used to evaluate their performance.
Considering the sensor technologies, metrological research has been conducted in terms of optical ToF (Xiong et al., 2017; Zanuttigh et al., 2016) and active stereo (Giancola et al., 2018; Wang and Shih, 2021). Furthermore, working principles of passive stereo sensors have been widely addressed in computer vision algorithms (Szeliski, 2022). Similar to our work, Chiu et al. (2019) and Halmetschlager-Funek (2019) jointly characterize ToF and active stereo.
With respect to the capture environments, related work examined the effects of object material (Xiong et al., 2017; Halmetschlager-Funek et al., 2019; Giancola et al., 2018; Hansard et al., 2012), color (Xiong et al., 2017; Giancola et al., 2018; Hansard et al., 2012), texture (Xiong et al., 2017; Hansard et al., 2012) and distance to objects (Halmetschlager-Funek et al., 2019). Furthermore, environmental lighting conditions (Halmetschlager-Funek et al., 2019) and multi-path effects (Giancola et al., 2018) were investigated. Specifically for ToF sensors, Wu et al. (2012) analyze multi-path effects originating from subsurface scattering and interreflections.
Moreover, we discuss related frameworks for jointly characterizing sensors. Halmetschlager-Funek et al. (2019) compare individually estimated depth values against manual measurements. Chiu et al. (2019) and Giancola et al. (2018) align the 3D data captured from sensors with real or synthetic ground truth data, respectively. Most similar to ours, Hansard et al. (2012) analyze ToF and structured light sensors using a spatial calibration and investigated object material, color, geometry, and texture using ground-truth data obtained from a structured light scanner.
2.2. Radio Wave Propagation and Range Sensing
So far, radio-frequency depth sensors (radar) were characterized in isolation. A more fundamental research direction examines the propagation of electromagnetic waves, which is the basis for RF ToF sensors. The general RF propagation behavior under different materials and geometries is measured by a parameter known as the radar cross section (RCS) (Knott et al., 2004). The RCS approximates the returned ratio of a transmitted radio signal and was measured in relation to a variety of materials (Knott et al., 2004; Semkin et al., 2020), as well as in the context of humans (Deep et al., 2020; Marchetti et al., 2018). Orthogonal research of Zhadobov et al. (2011) investigates the interaction of radio waves and human skin with respect to electromagnetic, thermal and biological aspects.
Moreover, studies of individual radar technologies have been conducted. Čopič Pucihar et al. (2022) evaluate the recognition of hand gestures using millimeter-wave radars in the presence of various materials. Wei et al. (2021) characterize imaging radars with respect to the geometry of metal objects in the context of security scanning. Furthermore, Bhutani et al. (2022) examine millimeter-wave radars at different frequencies, whereas Jha et al. (2018) analyze differences in their radiation between the near field and the far field. Sun et al. (2020) provide an overview of MIMO radars for autonomous driving, together with the characterization of their wave forms. Lastly, Ahmed (2021) presents millimeter-wave MIMO radar imaging systems in the context of security screening. To the best of our knowledge, no comprehensive characterization in conjunction with optical technologies has been done so far. Additionally, the existing efforts have been limited in scope with regard to RF depth sensing in the near field.
2.3. Fusion of RF and Optical Sensors in Close Range
Knowledge about complementary strengths is important for both, sensor characterization and sensor fusion. Research on multimodal sensor fusion in close range is very limited and mostly focused on capturing humans.
Zewge et al. (2019) perform people tracking with a $4\times 3$ MIMO radar and an active stereo camera. Similarly, Lee et al. (2023) propose a method for human pose estimation, which utilizes the data acquired from two $4\times 3$ MIMO radars synchronized with a monocular RGB camera. Both works do not utilize radar imaging methods due to the limited resolution. More similar to ours, Chen et al. (2023) use a high resolution $48\times 48$ MIMO radar and an RGB camera for human body reconstruction.
Furthermore, we address related datasets. Lim et al. (2021) introduce RaDICaL, an indoor and outdoor dataset of multiple people and objects, captured with a $4\times 3$ MIMO imaging radar and an active stereo camera. In the context of human body reconstruction, Chen et al. (2022) propose the mmBody benchmark that was captured with a $48\times 48$ MIMO radar and an RGB camera.
3. Preliminaries
As different research communities partially differ in their terminology, this paper pursues a unified terminology, summarized in the table below and used in the remainder of this paper.
[…]
mystyleleftmargin=0.0cm,rightmargin=0.0cm, innerleftmargin=0.1cm, innerrightmargin=0.1cm, innertopmargin=0.1cm, innerbottommargin=0.1cm {mdframed}[style=mystyle]Depth Imager. A sensor that, directly or indirectly, captures a depth image $\bm{D}$ of resolution $W\times H$, where each pixel $(u,v)$ contains a depth value $d$ measured along the axis perpendicular to the image plane. The depth may be indirectly measured from range and pixel position. We show the difference between range and depth in Figure 3.
Transmitter and Receiver. Optical receivers are small cells of image sensors, with a direct mapping to pixels. Transmitters are commonly LEDs or projectors. RF sensors have transmitting (TX) and receiving (RX) antennas.
Sensor. Describes all physical parts required for depth sensing and their spatial arrangement. Optical sensors typically consist of one or two cameras. Active sensors contain an additional illumination unit. RF sensors usually have one or more antenna arrays in different arrangements.
Depth Image Resolution $\mathbf{W\times H}$. The number of depth samples computed from the incoming signal. In cameras, the depth samples are directly computed for each pixel, i.e., each receiver. MIMO imaging radars apply signal post-processing to compute depth from the signal diversity at different receiver positions. Hence, the image resolution is not directly affected by the number of receivers but by the signal processing parameters such as the voxel density (cf. Section 4.2.2). While, in theory, the depth image resolution can be indefinitely high, in practice it is limited by the spatial resolution.
Spatial Resolution $\bm{\delta}$. The term resolution has several definitions. Here, we explicitly refer to spatial resolution as the minimum distance between two points in space that can be resolved from the received signal. Lower spatial resolution means higher minimum resolvable distance, so more incorrect measurements are made when points become inseparable, as seen in Figure 3. Spatial resolution is a theoretical measure, and external factors such as the sensor design can affect its effective resolution. Literature about traditional multi-static RF ToF sensors divides spatial resolution into range resolution ($\delta_{r}$), and cross-range resolution along the horizontal ($\delta_{h}$) and vertical ($\delta_{v}$) axes, respectively. While $\delta_{h}$ and $\delta_{v}$ are measured at a specific range, a more general formulation is the angular resolution, often referred to as azimuth ($\omega_{h}$) and elevation ($\omega_{v}$) resolution (Willis and Griffiths, 2007). Literature from the optical domain shares a similar definition, however, with a different terminology and refers to depth resolution (Zanuttigh et al., 2016) ($\delta_{z}$) and pixel resolution ($\delta_{h}$, $\delta_{v}$) instead. Similar to traditional RF ToF sensors, optical sensors assume that the target is situated in the far field, where depth is approximately the same as range and, hence, $\delta_{z}\approx\delta_{r}$. We note that this assumption does not generally hold in the near field. Due to the concept of an expanded antenna aperture with multiple transmitters and receivers, the resolution of near-field MIMO imaging radars is defined with respect to the three orthogonal axes $x,y,z$. In these, $z$ refers to the depth axis and $x,y$ are parallel to the antenna aperture. Contrary to their difference in definition, they share the same terminology as far-field RF ToF sensors such that $\delta_{z}$ is referred to as range resolution and $\delta_{x,y}$ is the cross-range (Ahmed, 2014) or lateral (Ahmed, 2021) resolution, respectively. We illustrate the respective definitions that are used for optical ($\delta_{h}$, $\delta_{v}$, $\delta_{r}$) and RF sensors ($\delta_{x}$, $\delta_{y}$, $\delta_{z}$) in Figure 3. For simplicity, we define spatial resolution only for the sensor center, where depth and range are approximately the same, such that $\delta_{z}\approx\delta_{r}$ and $\delta_{x,y}\approx\delta_{h,v}$.
Figure 3. Visualization of the effects caused by limited spatial resolution for multiple point targets. Optical sensors (left) have similar definitions as far-field RF ToF sensors and divide spatial resolution into depth $\delta_{r}\approx\delta_{z}$ and pixel resolution, $\delta_{h,v}$. Contrary to that, near-field imaging radars (right) refer to range $\delta_{z}$ and cross-range $\delta_{x,y}$ resolution. We define $\delta_{z}\approx\delta_{r}$ and $\delta_{x,y}\approx\delta_{h,v}$ for the sensor center, yet emphasize the conceptual difference between range and depth.
Refer to caption(Section 4.2.1)(Section 4.2.2)
Figure 4. Overview of the two depth sensing categories considered in this work. Spatially resolved methods compute the depth from disparity in the pixel positions. Time-resolved methods measure the depth through the round-trip propagation time of the received continuous wave (CW) signal. The types of wave forms utilized in our experiments are amplitude-modulated continuous wave (AMCW) and frequency-stepped continuous wave (FSCW).
4. Working Principles of Depth Imagers
In order to gain insight into the fundamental differences between optical and RF sensors, the first section characterizes wavelength-specific signal propagation. This is followed by an outline of the hardware design choices that are made for optical and RF depth imagers. After this, the working principles of the sensor technologies that are used in our experiments are discussed.
4.1. Characterization of Wavelength
Depth imagers are susceptible to the received, and optionally transmitted, signal wavelength. The wavelength affects both, the interaction of the signal with matter and the design of the sensor hardware and depth sensing algorithms. In this section we elaborate on both aspects, with a particular focus on the near-infrared light spectrum and the millimeter-wave (mmWave) radio-frequency spectrum.
NIR signals have a wavelength in the nanometer range. Given their high energy and strong interaction with matter, signal reflection or absorption is common, with scattering and non-diffractive phenomena dominating across most materials. Indirect effects on interactions with matter, therefore, often play a subordinate role, such that short propagation paths can be expected. Moreover, NIR light is pervasive in the environment, rendering optical technologies susceptible to external interference.
As suggested by their name, the wavelength of mmWave signals is longer by comparison. The low energy and reduced interaction with matter result in lower absorption and reflection, while there is a higher chance of a signal being transmitted through material. Specifically, the penetration depth of millimeter waves through matter is dependent on material parameters, such as, the resistivity and permittivity. For instance, the signals of security scanners can penetrate fabric but are primarily reflected on contact with metal objects (Ahmed, 2021). Furthermore, diffraction is more common with millimeter waves. This allows waves to bend around objects. Due to the aforementioned phenomena, the propagation paths of signals from active RF sensors are typically longer than of signals from optical sensors. Lastly, mmWave depth imagers operate with reduced external interference, as there are few natural microwave sources in the environment.
Due to the wavelength, the sensor design of RF sensors is inherently different from that of optical sensors. As stated by the general formulation of the Rayleigh criterion, the focus capacity and, hence, angular resolution $\omega$ of a sensor is limited by the signal wavelength $\lambda$ and the size of the sensor’s aperture $L$ (Hasch et al., 2012):
(1) $\omega_{x,y}=1.22\frac{\lambda}{L_{x,y}}\;.$
Optical sensors utilize camera lenses to refract the received signal, which enable a precise focus onto nanometer-sized pixels and, at the same time, exhibit a high angular resolution. In the context of the mmWave domain, a camera analogue can be conceptualized as a single-input multiple-output (SIMO) radar, which is to say, a sensor comprising a single transmitter and multiple receivers. As indicated by Equation 1, mmWave sensors have a significantly lower angular resolution than optical sensors. Thus, high-resolution SIMO radars require comparably large antenna arrays with large lenses, which has proven to be impractical. Instead, high-resolution RF imaging sensors often are synthetic aperture radars (SAR), which use digital beamforming to focus. They utilize the angle diversity from distinct transmitter and receiver positions, which form a virtual aperture of size $L$, to increase the angular resolution (Bliss and Forsythe, 2003) and require fewer antennas compared to SIMO systems. The majority of near-field SAR radars is implemented with MIMO arrays, that is, with multiple transmitters and receivers.
4.2. Depth Sensing Methods
In this section, we address the working principle of both optical and RF-based depth sensing methods used in our experiments. The content is organized in two categories: spatially resolved and time-resolved depth sensing, which are both depicted in Figure 4.
4.2.1. Spatially Resolved Depth Sensing
Spatially resolved depth imagers compute the point-wise depth from the respective pixel position in the image. In the following, we particularly address passive or active stereo(scopy) sensors.
Passive stereo sensors commonly utilize two cameras with a known relative spatial position to identify surface points in their respective images, a process known as correspondence or stereo matching (Szeliski, 2022). Given a correspondence pair of two pixels, the respective depth of this surface point is computed from their disparity. The quality of the correspondence matches affects the depth and accuracy of the results. Ambiguities in correspondence can arise due to textureless regions, poor lighting, motion or lens blur. Similarly, stereo matching can fail in terms of view-dependent effects or partial surface occlusions from one of the two receivers.
Active stereo sensors assist correspondence finding with an illumination unit that projects a pattern onto the target, usually in the NIR range, captured by the two cameras. The signal-multiplexed (Zanuttigh et al., 2016) pattern assists epipolar correspondence matching in addition to shading and texture cues, improving depth quality in textureless regions and low light. However, challenges include pattern distortions and signal oversaturation at the NIR receiver in bright conditions. Further details on spatially resolved depth sensors are provided in the supplementary material.
4.2.2. Time-Resolved Sensors (Time of Flight)
Time of Flight is an active depth sensing method, in which depth is derived from the round-trip propagation time that it takes for a signal to be transmitted and received. The majority of ToF sensors utilized in the near field employ continuous wave (CW) signal modulations, which measure time based on the relative phase shift $\Delta\varphi$ between the transmitted and received signal. The depth is derived from the range $r$, which is measured by (Zanuttigh et al., 2016):
(2) $r=\text{c}\frac{\Delta\varphi}{2\pi f}\;.$
The signal frequency is denoted as $f$, and c is the speed of light in vacuum, which closely matches that of light in air. For further details about the operating principle, we refer to the supplementary material. ToF technologies employ a simplified model for range sensing, which assumes that targets are diffuse scatterers, with each signal reflecting directly from the first target. As a result, these technologies are sensitive to multi-path interference, where signals propagate indirectly in multiple directions before reaching the receiver. In Section 8, we identify partially transmissive materials as a major cause of such interference. Furthermore, the unambiguous range of resolving $\Delta\varphi$ is limited to the periodicity of the sinusoidal CW signal. To extend this range, the carrier signal can be modulated over time. In our experiments, we use ToF sensors with amplitude-modulated continuous wave (AMCW) and frequency-stepped continuous wave (FSCW) signal modulations, which are illustrated in Figure 4. Up next, we will discuss the operating principles of these depth sensing methods.
NIR AMCW Time of Flight
AMCW ToF algorithms usually operate on the SIMO principle, as they do not require as expensive sensor apertures as imaging radars, and often have more receivers and transmitters than can be effectively managed computationally in MIMO depth estimation algorithms (Zanuttigh et al., 2016). They modulate the amplitude $A$ of a carrier signal over time $t$ using a repetitive modulation signal $s_{\text{m}}$ such that the transmitted signal $s_{\text{t}}$ can be described as:
(3) $s_{\text{t}}(t)=\underbrace{s_{\text{m}}(t)}_{\mbox{\small\raisebox{1.93748pt}{\smash{$A$}}}}\cdot\cos(2\pi tf+\phi_{\text{c}})\;.$
A constant phase offset is described by $\phi_{\text{c}}$. To extract the phase shift from the received signal $s_{\text{r}}$, it is cross-correlated with a so-called signal hypothesis $s_{\text{h}}$ (Zanuttigh et al., 2016):
[…]
$T_{\text{m}}$ is the period of the modulation signal $s_{\text{m}}$. Commonly, $s_{\text{h}}$ is chosen as the currently transmitted signal $s_{\text{t}}$ such that $c_{\text{r}}$ describes the signal similarity from which the relative phase shift to $s_{\text{r}}$ is inferred. Extracting this shift requires solving a multivariable equation system with parameters such as the received amplitude and external illumination. To achieve this, $c_{\text{r}}$ and, consequently, $s_{\text{r}}$ are commonly sampled at four points within $T_{\text{m}}$ (four-bucket-method) (Giancola et al., 2018). During the acquisition of those samples, AMCW ToF is affected by environmental changes, such as varying external NIR illumination and motion. Similar to active stereo, oversaturation of the NIR receiver may cause invalid signal responses.
MIMO FSCW Time of Flight
Given the frequency band $\text{b}=f_{\text{max}}-f_{\text{min}}$, FSCW ToF sensors iteratively send $\text{N}_{f}$ signals of one frequency in steps of $\Delta f=\text{b}/\text{N}_{f}$ (Bräunig et al., 2023b). More specifically, the transmitted signal $s_{\text{t}}$ of one capture can be described as:
(5) $s_{\text{t}}(t)$ $=A\cdot\cos(2\pi tf_{\text{m}}(n)+\phi_{\textit{c}})\text{ with }n=\left\lfloor t/\Delta t\right\rfloor$
(6) $f_{\text{m}}(n)$ $=f_{\text{min}}+\left(n\bmod{\text{N}_{f}}\right)\Delta f\;.$
We denote the time window of one frequency step as $\Delta t$. Time-division multiplexing (TDM) avoids signal interference and facilitates the separation of the received signal into its originating transmitter and frequency components. SAR signal processing computes the depth $d$ and the pixel position $(u,v)$ from the phase shift and the angular diversity originating from multiple transmitting and receiving positions. For MIMO imaging radars, the state-of-the-art algorithm of backprojection (BP) (Ahmed, 2021) computes confidence values about a target’s presence in 3D space. This is achieved on the basis of local feature distributions within a volume based on the integrated signal of each RX-TX antenna pair. Similar to the previous section, the BP algorithm computes a correlation between the received signal $s_{\text{r}}$ and a signal hypothesis $s_{\text{h}}$:
(7) $c_{\text{BP}}(\underbrace{x,y,z}_{\bm{p}})=\sum_{n=1}^{\text{N}_{f}}\sum_{i=1}^{\text{N}_{\text{RX}}}\sum_{j=1}^{\text{N}_{\text{TX}}}s_{\text{r}}(f_{n},\bm{r}_{i},\bm{t}_{j})s_{\text{h}}(f_{n},\bm{r}_{i},\bm{t}_{j},\bm{p})\;.$
A hypothesis is made on the basis of the transmitted signal, which is assumed to reflect at a point $\bm{p}\in\mathbb{R}^{3}$ in the sensor coordinate system, commonly sampled from a voxel grid of size $N_{v}=N_{x}\times N_{y}\times N_{z}$. The demodulated received signal, $s_{\text{r}}$, varies in transmit frequency $f_{n}=f_{\text{m}}(n)$, transmitter position $\bm{t}_{j}\in\mathbb{R}^{3}$, and receiver position $\bm{r}_{i}\in\mathbb{R}^{3}$. We further denote the numbers of transmitters and receivers as $\text{N}_{\text{TX}}$, and $\text{N}_{\text{RX}}$, respectively. Generally, hypotheses are made by assuming the signal propagation is following the so-called Born approximation (Ahmed, 2014). The result of the above equation is a complex phasor, $c_{\text{BP}}$, with its absolute value representing the confidence of a target’s presence, which is visualized in Figure 2. It is calculated from $s_{\text{r}}$ and $s_{\text{h}}$, which are analytic signals in complex notation. More details are given by Ahmed (2021). To compute a 2D depth map from the 3D voxel grid, an orthogonal maximum (intensity) projection (Bräunig et al., 2023b) is performed for each pixel $(u,v)=(x,y)$ along the cross-depth axes of the voxel grid:
[…]
Besides projection, the confidence values are used as thresholds for depth filtering, that is, to distinguish target depth from sidelobes and background noise. The absolute value of the confidence $c_{\text{BP}}$ depends on both, the received phase and amplitude. Besides the depth, reasons for varying amplitude and phase over different object materials and geometries are manifold, and further insights will be given in Section 9. As a result, it is challenging to generalize the depth filtering process for unknown objects. Similar to NIR AMCW ToF, a MIMO FSCW ToF sensor requires multiple signal samples to compute depth and, thus, is sensitive to environmental changes during capture.
ZED X Mini (2.2 mm) (Stereolabs, 2023) Realsense D435i (Intel, 2023) Azure Kinect (Microsoft, 2022) QAR5 (Submodule) (Rohde & Schwarz, 2023)
Manufacturer Stereolabs Intel Microsoft Rohde & Schwarz
Depth Sensing Technology Passive Stereoscopy Active Stereoscopy Time of Flight (NIR) Time of Flight (RF)
[…]
Capture Frame Rate 30 fps 30 fps 30 fps $\approx$ 70 fps*
Transmitters Type -- Laser Projector LED Array TDM Antenna Array
Array Size -- -- -- 2$\times$47 TX $\updownarrow$
Wavelength -- 840–860 nm 860 nm 3.6-4.2 mm
Frequency -- $\approx$ 353 THz $\approx$ 353 THz 72-82 GHz
Signal Modulation -- Spatial Multiplexing AMCW FSCW ($\text{N}_{f}=128$)
Receivers Type Image sensor Image sensor Image Sensor Antenna Array
Array Size 2$\times$1928$\times$1208 px 2$\times$1280$\times$800 px 1024$\times$1024 px 2$\times$47 RX $\leftrightarrow$
Spatial Size* 2$\times$5.8$\times$3.6 mm 2$\times$3.8$\times$2.4 mm 3.6$\times$3.6 mm 2$\times$13.8 cm
Depth Image Resolution 1920$\times$1080 px 1280$\times$720 px 640$\times$576 px 301$\times$301 px
[…]
Table 1. Overview of the sensors and their parameters used in our experiments. Rows with * indicate information not directly given by the manufacturer. $\delta_{x}$ and $\delta_{y}$ of camera-based systems is approximately determined from the per-pixel field of view. Spatial resolution formulae are provided in the supplementary material. Due to missing data for the Azure Kinect, $\delta_{z}$ is assumed to be theoretically higher than the depth accuracy given in (Bamji et al., 2018).
5. The MAROON Dataset
The capture of the MAROON dataset allows for a comprehensive analysis with respect to the characteristics of the four previously described depth sensing techniques. This section outlines the capture setup and data acquisition pipeline, depicted in Figure 1, to aid future research on the publicly accessible data. Four single-view depth sensors are used in our experiments: Stereolabs ZED X (Stereolabs, 2023) (Passive Stereo), Intel Realsense D435i (Intel, 2023) (Active Stereo), Microsoft Azure Kinect (Microsoft, 2022) (NIR ToF), and a submodule of Rohde & Schwarz’s QAR50 (Rohde & Schwarz, 2023) (RF ToF). Example data captured from these sensors is visualized in Figure 2.
5.1. Sensor Setup
Our sensor setup consists of four mounted single-view depth sensors and a ground truth (GT) optical multi-view stereo (MVS) system comprising five calibrated DSLR cameras, which are depicted on the left in Figure 1. While all single-view sensors are designed to achieve an optimal balance between depth quality and acquisition time, the MVS system employs an offline reconstruction process that is specifically optimized for depth quality. In summary, eight cameras are mounted on tripods and arranged around the MIMO imaging radar on a desk, thereby maximizing the area of intersection of each sensor’s field of view, to ensure similar object visibility. All sensors and the GT system are time-synchronized, either through hardware or software, to capture the object at the same moment.
Prior to capture, the object is positioned in the center of the squared radar aperture, and approximately at the center of the joint field-of-view intersection, propped up with boards crafted from styrofoam — a material that is considered to be nearly fully penetrated by the RF signal — to prevent external interference of radio waves from other sources in the vicinity, apart from the object of interest. For similar reasons, absorbers are placed behind the object of interest. Similarly, for optical sensors a loose black cloth, which is penetrated by RF signals, is suspended in front of the absorbers to visually occlude the room’s background. The sensor settings are chosen with respect to a trade-off between fair sensor comparability and practical applicability. More details can be found in the supplementary material. An overview of the chosen settings, together with relevant sensor parameters, is given in Table 1.
5.2. Data Acquisition Pipeline
To examine the characteristics of various objects in relation to different materials, geometries, and distances from the sensor, we recorded the MAROON dataset, which comprises static and quasi-static targets captured from all sensors simultaneously. With respect to the order of steps described in Figure 1 (middle), we will now continue to elaborate on the details of the acquisition pipeline.
Spatial Calibration
We spatially aligned the coordinate systems of each depth imager using the calibration method by Wirth et al. (2024). In this method, four respective spherical objects of styrofoam and metal, tailored to the visibility of optical and RF sensors, are captured. In the sensor-specific reconstructions, these spheres are automatically located and jointly aligned using spatial registration. This approach enables a direct comparison of the object reconstructions in a metrical space. Calibration errors are expected to be in 1–2 mm range with respect to the Chamfer distance, in analogy with Wirth et al. (2024).
The five DSLR cameras of the MVS system are treated as a unified sensor with a common coordinate system, which is spatially calibrated with that of the depth imagers. The camera extrinsic and intrinsic parameters of the MVS system are determined from images capturing a conventional optical calibration target with a checkerboard pattern. For this, we use the software provided by Agisoft Metashape. Remaining calibration errors exhibit a root mean square reprojection error of 0.38 px, averaged over all camera calibrations performed during the dataset capture.
Object Preparation and Capture
The reconstruction method of MVS is similar to passive stereo imagers. Hence, inaccurate reconstructions can be the result when dealing with textureless and view-dependent object materials. To circumvent this limitation, we implement a distinct capture process for particularly challenging objects to generate more reliable GT reconstructions. Once the object has been captured by all depth imagers, a thin multicolored speckle pattern is applied using water colors that assists the correspondence finding by a subsequent MVS-only capture. In order to ensure exact alignment between that GT reconstruction and other imaging modalities, the speckle is applied in situ without moving the object.
In total, each object is recorded at three different distances to the MIMO imaging radar of 30 cm, 40 cm, and 50 cm, respectively. The remaining depth imagers are situated behind the radar. Their corresponding object-to-sensor distance is determined from the distance to the radar and from the relative position between each optical sensor and the imaging radar, which is given by the calibration parameters. Based on the Euclidean norm of the mean translation across all calibrations conducted, we report an additional object-to-sensor distance of +8 cm (Azure Kinect), +6 cm (Realsense D435i), and +5 cm (ZED X Mini), respectively. We record 20 frames for each optical sensor and 10 radar frames. In total, we capture 45 objects and list further statistics about the dataset in Table 2.
Optical Segmentation
To perform an accurate and precise object-centric sensor evaluation, it is essential to isolate the estimated object depth from the background. For optical systems, we acquire segmentation masks by performing a semi-automatic foreground-background segmentation. Given that all depth imagers capture RGB images — either for depth estimation or via a separate calibrated camera — we first segment the RGB images using manually defined object labels in conjunction with Grounded-SAM (Ren et al., 2024). This generates a binary segmentation mask of the object, $\bm{M}$, where all valid pixels $(u,v)$ are included in $\bm{M}^{+}(u,v)=\{\bm{M}(u,v)>0\}$. We then manually correct failure cases in the resulting segmentation masks. The same procedure is employed to MVS images to produce masked GT reconstructions. For the imaging radar, the voxel volume of the BP algorithm (Equation 7) is constrained to enclose only the object of interest. In this way, segmentation masks are automatically determined from the valid pixels remaining after depth estimation.
Depth Estimation
[…]
Statistics MAROON
# objects 45
# static objects 41
# quasi-static objects 4
# prepared speckled objects 14
# captures (# objects $\times$ 3 distances) 135
# total / unique optical depth frames 8100 / 405
# total / unique RF depth frames 1350 / 135
Table 2. Statistics of the MAROON dataset. Assuming that all captured objects are static, the number of total frames include duplicate captures, possibly varying in random depth noise, while the unique frames only contain one capture per object of each sensor.
6. Evaluation
In this section, we compare the reconstructions produced by the four presented depth imagers with a ground-truth reconstruction in a common metric space and describe the methods used in this process. Subsequently, the results of these methods are presented.
6.1. Methods
In accordance with the order in Figure 1 (right), the necessary steps for evaluation are explained.
Reconstruction
We acquire a point cloud of the object’s surface utilizing the 2D depth and auxiliary data provided by MAROON. For a given pixel position $(u,v)$ and its corresponding depth $d$ from an optical depth sensor, we first verify its validity using the segmentation map of the same resolution — a step that has already been performed for radar during depth filtering. Subsequently we project each valid triple $(u,v,d)$ back into 3D space using the given transformation matrix $\bm{T}\in\mathbb{R}^{4\times 4}$:
[…]
For all optical depth imagers, this equation is the inverse of a perspective transformation with intrinsic camera matrix $\bm{I}\in\mathbb{R}^{3\times 3}$, pixel offset vector $\bm{t}\in\mathbb{R}^{3}=\bm{0}$ and $a=d$. Analogously for radar data, the equation is the inverse of an orthographic transformation with a scale matrix $\bm{I}$, pixel offset $\bm{t}$, and $a=1$.
During evaluation, we average valid depth values across 10 frames for static objects to incorporate temporal characteristics and reduce random noise. We do not average quasi-static objects (e.g., human hands) and instead take the first frame, as it is closest to the point in time where the GT captures have been taken.
Joint Alignment
To estimate the deviation of a sensor reconstruction $\bm{R}_{s}\in\mathbb{R}^{M\times 3}$ from the GT, $\bm{R}_{g}\in\mathbb{R}^{N\times 3}$, we use the previously determined spatial calibration parameters ${\bm{K}_{g\rightarrow s}\in\mathbb{R}^{4\times 4}}$ to transform $\bm{R}_{g}$ from the GT space $g$ into the sensor space $s$:
(10) $\widetilde{\bm{R}}_{g}^{s}=\widetilde{\bm{R}}_{g}\bm{K}_{g\rightarrow s}^{T}\;.$
$\widetilde{\bm{R}}$ denotes the homogeneous version of $\bm{R}$. We use the notation $\bm{R}^{*}$ to indicate a reconstruction that has been transformed to sensor space *.
Depth Deviation Measurement
For each object, we compute the point-wise deviation between a sensor and the GT reconstruction with respect to two metrics: one-sided Chamfer distance and one-sided projective error. The one-sided Chamfer distance, C, is computed per point $\bm{p}\in\mathbb{R}^{3}$ in the source point cloud $\bm{P}\in\mathbb{R}^{N\times 3}$ with respect to the distance to the nearest point $\bm{q}\in\mathbb{R}^{3}$ in the destination point cloud $\bm{Q}\in\mathbb{R}^{M\times 3}$:
(11) $\text{C}_{\bm{p}}(\bm{Q})=\underset{\bm{q}\in\bm{Q}}{\text{min}}\lVert\bm{p}-\bm{q}\lVert_{2}\;.$
The one-sided projective depth error P is computed per pixel $(u,v)$ of two depth maps $\bm{D}\in\mathbb{R}^{W\times H}$ and $\bm{F}\in\mathbb{R}^{W\times H}$ of a common image plane:
(12) $\text{P}_{u,v}(\bm{D},\bm{F})=\lvert\bm{D}(u,v)-\bm{F}(u,v)\rvert\;.$
For both, C and P, the respective subscripts $p$ and $u,v$ are used as placeholders to denote the points and pixels used in the metric computation. We compute both metric types for each reconstruction pair of a sensor, $\bm{R}_{s}$, and the GT, $\bm{R}_{g}$. Since the one-sided Chamfer distances are sensitive to the point cloud density, we uniformly sample the points from the sensor and the GT with respect to a common image pixel grid. We achieve this by, first, computing a simulated depth map $\widehat{\bm{D}}_{g}^{s}$ in the image space of the sensor $s$ and, second, reconstructing $\widehat{\bm{R}}^{s}_{g}$ from this depth map, using Equation 9 with the sensor-specific transformation $\bm{T}_{s}$. The simulated depth map is computed by rasterizing a triangulated representation of $\bm{R}^{s}_{g}$ with respect to $\bm{T}_{s}$. In this way, we also discard points in optical sensors $\bm{R}^{s}_{g}$ that are not visible in the view of camera $s$. The resulting depth map $\widehat{\bm{D}}_{g}^{s}$ is additionally used to measure the projective error. To summarize, we compute: $\text{C}_{\text{g}}$ Chamfer distance ground truth. $\forall\bm{g}\in\widehat{\bm{R}}^{s}_{g}:\text{C}_{\bm{g}}(\bm{R}_{s})$ $\text{C}_{\text{s}}$ Chamfer distance sensor. $\forall\bm{s}\in\bm{R}_{s}:\text{C}_{\bm{s}}(\widehat{\bm{R}}^{s}_{g})$ P Projective error. $\forall(u,v)\in\bm{M}^{+}(u,v):\text{P}_{u,v}(\bm{D}_{s},\widehat{\bm{D}}_{g}^{s})$ for ${\bm{M}=\bm{M}_{s}\cap\widehat{\bm{M}}_{g}^{s}}$, where $\bm{M}_{s}=\{\bm{D}_{s}>0\}$ and $\widehat{\bm{M}}_{g}^{s}=\{\widehat{\bm{D}}_{g}^{s}>0\}$ describe the intersection masks of valid pixels from the sensor and the projected GT, respectively. $\text{P}_{\text{e}}$ Projective error with erosion. $\forall(u,v)\in\bm{M}_{e}^{+}(u,v):\text{P}_{u,v}(\bm{D}_{s},\widehat{\bm{D}}_{g}^{s})$ for ${\bm{M}_{e}=\bm{M}_{s}\cap f(\widehat{\bm{M}}_{g}^{s})}$, where $f(\widehat{\bm{M}}_{g}^{s})$ is a function performing mask erosion using a kernel size of $K\times K$ pixels. The size $K\in[0,20]$ is semi-manually selected for each object and sensor, and included together with other evaluation metadata in the release of our dataset.
Metric Type Silhouette Noise Missing Surfaces 3D Error Depth Error
$\text{C}_{\text{g}}$  —  —
$\text{C}_{\text{s}}$  —  —
P  —  —
$\text{P}_{\text{e}}$  —  —  —
Table 3. Categorization of the presented metrics with respect to their sensitivities. In addition to depth, a 3D error evaluates errors along the cross-depth axes.
6.2. Results
In this section, we first present the evaluation results, quantified using four complementary metrics: $\text{C}_{\text{g}}$, $\text{C}_{\text{s}}$, P, and $\text{P}_{\text{e}}$. Each metric is sensitive to different aspects, as detailed in Table 3. Important to note is that the computed quantities do not perfectly reflect the absolute depth error in the real world, due to the limitations of our experimental setup, which is constrained by the accuracy of the GT system and the spatial calibration. Instead, the results should be interpreted as depth deviation from the ground truth, referred to as sensor-GT deviation. We begin by presenting the sensor-GT deviation results in relation to various objects and different object-to-sensor distances. Given that near-field imaging radars are less explored compared to optical depth sensors, we dedicate the latter part of this section to the radio-frequency signal response. Finally, we note that in this section the results are objectively presented, reserving further interpretations for subsequent sections, where they will be discussed from a general perspective (Section 7), with specific attention to partially transmissive media (Section 8), and focusing on the RF signal response (Section 9).
6.2.1. Depth Deviation at Single and Multiple Object Distances
In Table 4, we list the mean $\mu$ and standard deviation $\sigma$ of each metric type with respect to 12 selected objects from the MAROON dataset. These objects were positioned at an object-to-sensor distance of 30 cm. We provide object images and a comprehensive evaluation of all 45 objects in the supplementary material. For completeness, we give a brief overview of the overall statistics by investigating the number of best and worst results across all objects for $\mu$, respectively:
  • RF ToF: performs worst in $\text{C}_{\text{g}}$
  • NIR ToF: performs worst in $\text{C}_{\text{s}}$, P, $\text{P}_{\text{e}}$
  • Active Stereo: performs best in $\text{C}_{\text{g}}$, $\text{C}_{\text{s}}$, P, $\text{P}_{\text{e}}$
  • Passive Stereo: performs neither best or worst
To give an intuition on relative depth deviations between sensors, we present the median, mean, and standard deviation, denoted as $\widetilde{\mu}/\mu~{}(\pm\sigma)$, calculated across the differences in metric values for all pairwise sensor combinations. The results for each metric type are:
  • $\text{C}_{\text{g}}$$0.23~{}\text{cm}/0.48~{}\text{cm}~{}(\pm 0.61~{}\text{cm})$
  • $\text{C}_{\text{s}}$$0.19~{}\text{cm}/1.06~{}\text{cm}~{}(\pm 3.53~{}\text{cm})$
  • P: $0.34~{}\text{cm}/1.58~{}\text{cm}~{}(\pm 4.36~{}\text{cm})$
  • $\text{P}_{\text{e}}$$0.31~{}\text{cm}/1.78~{}\text{cm}~{}(\pm 5.14~{}\text{cm})$
Additionally, we illustrate the distribution of the depth deviation across all objects for varying placement distances of 30 cm, 40 cm, and 50 cm. The results are displayed as boxplots in Figure 5.
Metric Type Cardboard Sponge Scrubber Plushie Tape Dispenser Statue
RF ToF $\text{C}_{\text{g}}$ 0.13 ($\pm$ 0.06) 1.52 ($\pm$ 0.97) 0.58 ($\pm$ 0.29) 0.81 ($\pm$ 0.47) 0.31 ($\pm$ 0.23) 0.27 ($\pm$ 0.25)
NIR ToF 0.10 ($\pm$ 0.06) 0.79 ($\pm$ 0.45) 0.64 ($\pm$ 0.34) 0.49 ($\pm$ 0.21) 0.84 ($\pm$ 0.30) 0.32 ($\pm$ 0.28)
Active Stereo 0.08 ($\pm$ 0.05) 0.18 ($\pm$ 0.17) 0.20 ($\pm$ 0.16) 0.13 ($\pm$ 0.14) 0.16 ($\pm$ 0.14) 0.16 ($\pm$ 0.13)
Passive Stereo 0.24 ($\pm$ 0.11) 0.26 ($\pm$ 0.15) 0.14 ($\pm$ 0.11) 0.19 ($\pm$ 0.21) 0.15 ($\pm$ 0.11) 0.13 ($\pm$ 0.13)
RF ToF $\text{C}_{\text{s}}$ 0.15 ($\pm$ 0.08) 0.54 ($\pm$ 0.40) 0.97 ($\pm$ 0.61) 2.21 ($\pm$ 2.05) 0.55 ($\pm$ 0.77) 0.17 ($\pm$ 0.11)
NIR ToF 0.16 ($\pm$ 0.18) 0.79 ($\pm$ 0.42) 0.59 ($\pm$ 0.30) 0.53 ($\pm$ 0.28) 1.16 ($\pm$ 0.60) 0.43 ($\pm$ 0.42)
Active Stereo 0.08 ($\pm$ 0.05) 0.17 ($\pm$ 0.15) 0.17 ($\pm$ 0.11) 0.12 ($\pm$ 0.38) 0.17 ($\pm$ 0.17) 0.19 ($\pm$ 0.76)
Passive Stereo 0.30 ($\pm$ 0.13) 0.33 ($\pm$ 0.18) 0.19 ($\pm$ 0.13) 0.23 ($\pm$ 0.27) 0.22 ($\pm$ 0.20) 0.18 ($\pm$ 0.69)
RF ToF P 0.13 ($\pm$ 0.14) 2.73 ($\pm$ 2.47) 1.28 ($\pm$ 0.76) 3.64 ($\pm$ 3.23) 0.67 ($\pm$ 1.08) 0.20 ($\pm$ 0.26)
NIR ToF 0.19 ($\pm$ 0.25) 1.32 ($\pm$ 0.58) 0.91 ($\pm$ 0.47) 0.85 ($\pm$ 0.52) 1.57 ($\pm$ 0.70) 0.77 ($\pm$ 3.13)
Active Stereo 0.10 ($\pm$ 0.12) 0.29 ($\pm$ 0.42) 0.27 ($\pm$ 0.34) 0.24 ($\pm$ 1.25) 0.24 ($\pm$ 0.35) 0.90 ($\pm$ 5.17)
Passive Stereo 0.36 ($\pm$ 0.17) 0.47 ($\pm$ 0.51) 0.29 ($\pm$ 0.38) 0.35 ($\pm$ 0.54) 0.29 ($\pm$ 0.35) 1.43 ($\pm$ 7.09)
RF ToF $\text{P}_{\text{e}}$ 0.12 ($\pm$ 0.13) 2.93 ($\pm$ 2.60) 1.35 ($\pm$ 0.66) 3.61 ($\pm$ 3.23) 0.66 ($\pm$ 1.07) 0.20 ($\pm$ 0.27)
NIR ToF 0.08 ($\pm$ 0.10) 1.69 ($\pm$ 0.23) 1.16 ($\pm$ 0.27) 0.82 ($\pm$ 0.38) 1.70 ($\pm$ 0.89) 0.25 ($\pm$ 0.13)
Active Stereo 0.06 ($\pm$ 0.08) 0.42 ($\pm$ 0.50) 0.22 ($\pm$ 0.27) 0.16 ($\pm$ 0.30) 0.15 ($\pm$ 0.21) 0.16 ($\pm$ 0.21)
Passive Stereo 0.38 ($\pm$ 0.14) 0.55 ($\pm$ 0.56) 0.24 ($\pm$ 0.29) 0.18 ($\pm$ 0.26) 0.20 ($\pm$ 0.22) 0.10 ($\pm$ 0.13)
Metric Type S1 Hand Open Hand Printed Flat Mirror Candle Flowerpot (Transparent) V1 Metal Plate
RF ToF $\text{C}_{\text{g}}$ 0.36 ($\pm$ 0.38) 0.71 ($\pm$ 0.78) 0.87 ($\pm$ 0.26) 1.50 ($\pm$ 1.12) 1.31 ($\pm$ 1.21) 0.12 ($\pm$ 0.05)
NIR ToF 0.31 ($\pm$ 0.14) 0.25 ($\pm$ 0.12) 3.77 ($\pm$ 1.97) 2.04 ($\pm$ 0.40) 2.73 ($\pm$ 1.03) 0.77 ($\pm$ 0.42)
Active Stereo 0.12 ($\pm$ 0.09) 0.09 ($\pm$ 0.07) 2.13 ($\pm$ 1.52) 0.26 ($\pm$ 0.29) 0.74 ($\pm$ 0.53) 0.08 ($\pm$ 0.06)
Passive Stereo 0.20 ($\pm$ 0.16) 0.21 ($\pm$ 0.37) 2.31 ($\pm$ 1.61) 1.64 ($\pm$ 0.78) 2.01 ($\pm$ 0.83) 0.13 ($\pm$ 0.07)
RF ToF $\text{C}_{\text{s}}$ 0.22 ($\pm$ 0.15) 0.17 ($\pm$ 0.13) 0.91 ($\pm$ 0.14) 5.57 ($\pm$ 2.78) 1.86 ($\pm$ 2.41) 0.13 ($\pm$ 0.06)
NIR ToF 0.38 ($\pm$ 0.26) 0.29 ($\pm$ 0.20) 33.31 ($\pm$ 9.07) 1.71 ($\pm$ 0.49) 3.10 ($\pm$ 1.22) 0.81 ($\pm$ 0.43)
Active Stereo 0.13 ($\pm$ 0.10) 0.09 ($\pm$ 0.06) 30.21 ($\pm$ 14.59) 0.25 ($\pm$ 0.26) 1.27 ($\pm$ 1.78) 0.09 ($\pm$ 0.07)
Passive Stereo 0.26 ($\pm$ 0.22) 0.18 ($\pm$ 0.34) 27.02 ($\pm$ 11.33) 1.28 ($\pm$ 0.65) 1.86 ($\pm$ 0.93) 0.16 ($\pm$ 0.11)
RF ToF P 0.22 ($\pm$ 0.25) 0.16 ($\pm$ 0.20) 0.93 ($\pm$ 0.12) 7.41 ($\pm$ 3.79) 2.74 ($\pm$ 3.66) 0.11 ($\pm$ 0.12)
NIR ToF 0.52 ($\pm$ 0.43) 0.33 ($\pm$ 0.29) 37.84 ($\pm$ 14.84) 2.78 ($\pm$ 0.35) 5.24 ($\pm$ 2.04) 0.95 ($\pm$ 0.48)
Active Stereo 0.22 ($\pm$ 1.25) 0.16 ($\pm$ 1.30) 39.66 ($\pm$ 24.75) 0.42 ($\pm$ 0.49) 2.08 ($\pm$ 2.30) 0.10 ($\pm$ 0.13)
Passive Stereo 0.35 ($\pm$ 0.41) 1.73 ($\pm$ 8.95) 30.82 ($\pm$ 14.01) 2.10 ($\pm$ 0.98) 3.50 ($\pm$ 1.37) 0.19 ($\pm$ 0.15)
RF ToF $\text{P}_{\text{e}}$ 0.22 ($\pm$ 0.25) 0.16 ($\pm$ 0.20) 0.93 ($\pm$ 0.12) 7.37 ($\pm$ 3.85) 2.76 ($\pm$ 3.66) 0.10 ($\pm$ 0.11)
NIR ToF 0.51 ($\pm$ 0.27) 0.30 ($\pm$ 0.09) 39.68 ($\pm$ 6.57) 2.75 ($\pm$ 0.15) 6.18 ($\pm$ 1.79) 0.79 ($\pm$ 0.39)
Active Stereo 0.16 ($\pm$ 0.24) 0.08 ($\pm$ 0.10) 43.84 ($\pm$ 20.28) 0.31 ($\pm$ 0.44) 2.52 ($\pm$ 1.77) 0.07 ($\pm$ 0.09)
Passive Stereo 0.25 ($\pm$ 0.34) 0.17 ($\pm$ 0.15) 35.96 ($\pm$ 7.83) 2.15 ($\pm$ 0.65) 4.36 ($\pm$ 0.71) 0.15 ($\pm$ 0.09)
Table 4. We measure the sensor-GT deviation with respect to $\text{C}_{\text{g}}$, $\text{C}_{\text{s}}$, P, and $\text{P}_{\text{e}}$, which we list in the form $(\mu\pm\sigma)$, consisting of the mean $\mu$ and standard deviation $\sigma$ in centimeters, computed over the entire metric domain, respectively. The best results among all sensors of one metric type are highlighted in bold and the worst results are underlined. The results are discussed in Section 7.
Figure 5. We plot the distribution of the mean error across all objects with respect to different object-to-sensor distances. Solid (—) and dashed (- -) horizontal lines indicate the median and the mean of the distribution, respectively. The results are discussed in Section 7.
Figure 6. We present the received signal magnitude (mean absolute phasor value) in relation to the mean depth deviation of metric $\text{C}_{\text{g}}$ and further investigate both quantities with respect to varying object material, geometry (median surface incidence angle), and size (relative surface area). We highlight measurements where large objects appear outside the radar’s field of view (FOV) in grey regions, as they exhibit comparatively higher depth deviations compared to the ground truth reconstructions that possibly extend beyond this FOV. The results are discussed in Section 9.
6.2.2. RF Signal Response at Single Object Distance
In addition to the phase shift, the point-wise confidence of the backprojection algorithm for RF ToF is significantly affected by the signal amplitude; however, disentangling amplitude and phase in the presence of signal interference imposes the same challenges that arise from recovering the phase shift itself. To avoid inducing additional bias through the assumptions made in signal processing, we investigate the unprocessed signal response in the MIMO imaging radar across multiple objects. For each object at 30 cm object-to-sensor distance, we compute the mean absolute value out of all complex phasors received from the raw signal of size $\text{N}_{\text{RX}}\times\text{N}_{\text{TX}}\times N_{f}$, averaged over 10 frames. We refer to this quantity as signal magnitude, emphasizing the difference from signal amplitude.
To investigate correlations between signal response and reconstruction quality, we analyzed the signal magnitude relative to the depth deviation from ground truth, expressed as the mean of $\text{C}_{\text{g}}$. These findings are shown on the left in Figure 6. We further differentiate between signal magnitude (upper row) and depth deviation (bottom row), examining their relationships with respect to object material, geometry, and size from left to right. Object materials are categorized into six classes, with detailed information available in the supplementary material. The goal of this classification is to highlight material differences on a coarse level, noting the large object variety that still persists within one material class. The object geometry is quantified by the median angle in degrees between the point-wise surface normals of the GT reconstruction and the depth direction (along the $z$-axis) of the imaging radar. As we positioned the objects to align their primary orientation with the viewing direction of the planar square-shaped antenna aperture — which particularly becomes important for flat objects — the median angle mainly reflects geometric complexity, with objects having a higher surface incidence angle showing larger portions oriented away. Object size is determined by the relative surface area compared to the radar antenna aperture. It is computed from the fraction of the object’s 2D axis-aligned bounding box $A$ (in the $x$- and $y$-axis) inside of the 2D axis-aligned bounding box $B$ of the antenna array by using the formula: $(A\cap B)/B$.
7. Sensor Characterization by Depth Deviation
First, we present complementary perspectives on the data, where we put the depth deviation of all 45 objects into relation with the different metric types and, subsequently, the different types of depth imagers. We analyze each representation in turn, highlighting common trends, in combination with previously stated results of Section 6.2. Finally, we examine the depth deviation obtained specifically for the selected objects of Table 4.
7.1. General Trends
Interpreting the extensive numerical data on depth deviations in Table 4 and the supplementary material can be challenging, so we provide visual, complementary views in this section. In order to relate different quantities to each other, we use barycentric interpolation based on triples of metric types (Figure 7) and sensors (Figure 8), respectively. For each triple $(\mu_{a},\mu_{b},\mu_{c})$, the mean values for depth deviation ($\mu$) of each object are transformed to affine coordinates $(w_{a},w_{b},w_{c})$ by using the formula $w_{\{a,b,c\}}=\mu_{\{a,b,c\}}/(\mu_{a}+\mu_{b}+\mu_{c})$. Circle locations closer to a triangle corner indicate higher relative depth deviation. Moreover, as the triples are drawn from a set of four, the triangles are arranged in the shape of an unfolded tetrahedron, highlighting that each triangle’s contents can be seen as a projection of barycentric coordinates within a (3D) tetrahedron $w_{\{a,b,c,d\}}=w_{\{a,b,c,d\}}/(w_{a}+\cdots+w_{d})$.
Figure 7. A complementary view on the depth deviation across different metric types. For each triplet of metrics $M_{i}\in\{\text{$\text{C}_{\text{g}}$},\text{$\text{C}_{\text{s}}$},\text{P},\text{$\text{P}_{\text{e}}$}\}$, we convert each mean depth deviation $\mu_{i}$ to affine coordinates, $\bar{\mu}_{i}=\mu_{i}/\sum_{i}\mu_{i}$, that map an object’s errors in to a triangle whose corners correspond to metrics $M_{i}$. All 45 MAROON objects are shown as circles, with selected objects from Table 4 highlighted in solid colors. Samples closer to a triangle corner indicate a higher relative depth deviation in the corresponding metric.
Figure 8. A complementary view on the depth deviation across different sensors. The sensors are denoted as R (RF ToF), N (NIR ToF), A (Active Stereo), and P (Passive Stereo), respectively. Analogously to Figure 7, we convert the mean depth deviations $\mu_{i}$ to affine coordinates within triangles corresponding to all possible sensor triples. All 45 MAROON objects are shown as circles, with selected objects from Table 4 highlighted in solid colors. Samples closer to a triangle corner indicate a higher relative depth deviation for the corresponding sensor.
In Figure 7, we provide a qualitative comparison of each sensor’s depth deviation from GT with respect to the four presented metrics. In dense reconstructions, as is typical for optical depth sensors, metrics based on nearest neigbors (here, $\text{C}_{\text{g}}$ and $\text{C}_{\text{s}}$) are bound to be lower than those based on projection (P and $\text{P}_{\text{e}}$); they also tend to be more resilient against noise. For RF reconstructions, however, that are prone to sparse depth maps, Chamfer distances often create false matches; accordingly, $\text{C}_{\text{g}}$ dominates for RF ToF compared to the other metrics. For further discussion regarding the sparsity of RF reconstructions, see Section 9.
For optical sensors, a marginal trend towards the corners of $\text{C}_{\text{s}}$ and P, away from the silhouette-resilient $\text{P}_{\text{e}}$ and $\text{C}_{\text{g}}$, cf. Table 3, indicates the presence of noise at object silhouettes.
In Figure 8, we observe a considerable spread of depth deviations across different sensors. As noted in Section 6.2, the relative depth deviation between sensors ranges from 1.9 to 3.4 mm, that is, the variation in the depicted normalized error occurs within a comparatively small range of absolute errors. As a general trend, the two stereo sensors have the lowest depth deviation (see triangles N–R–A and N–R–W), with a moderate edge for active stereo, particularly for metrics $P$ and $P_{\text{e}}$, see the $A$$P$ axis, where passive stereo‘s performance degrades. This is consistent with the results in Section 6.2.1, where we find that the active stereo has the highest number of best results. However, given that relative depth deviations between sensors differ by only a few millimeters, we conclude that the seemingly improved depth quality of active stereo sensors is of minimal significance. Moreover, an examination of the scatter plot reveals a marginal weight trend towards the corners of the two ToF sensors. Reconstruction errors between the two ToF sensors are highly object-dependent, and many objects, including many of the highlighted ones, demonstrate multi-path effects due to perfect signal reflections, retro-reflectivity, or partial signal transmission. Further details on this will be discussed in the next section.
Considering the results over varying object-to-sensor distances in Figure 5, we observe that the depth deviation of the passive stereo and the NIR ToF sensor is significantly more distance-specific compared to active stereo and RF ToF. For passive stereo, this may be due to a decrease in effective spatial resolution, where $\delta_{z}$ directly depends on $\delta_{x,y}$ (see supplementary). As the distance between the object and the sensor increases, $\delta_{x,y}$ decreases, resulting in a loss of high-frequency color details while the object appears smaller in the image. Compared to passive stereo, the active stereo sensor has a comparably higher effective resolution, assuming that the resolution of both sensors differs in accordance with $\delta_{xyz}$ in Table 1. The unique active NIR pattern may also be less sensitive to decreases in spatial resolution, maintaining the quality of correspondence matches. The trend for the NIR ToF sensor aligns with findings from Bamji et al. (2018) for 30–50 cm distances. However, we argue that absolute errors for a target with 20% reflectivity do not fully represent all objects in our experiments. We suggest that, in addition to the expected decrease in spatial resolution, a greater depth deviation with increasing distance arises from the signal-to-noise ratio with respect to environmental light, which typically decreases over distance due to the inverse-square law.
Notably, in our results, RF ToF does not exhibit a distance-dependent depth deviation, unlike its optical counterpart. This seems to contradict Table 1, where $\delta_{z}$ for RF ToF degrades more rapidly with depth, compared to optical sensors; however, the theoretical decay with depth stems from the worsening separability of neighboring point targets, which is not a pertinent scenario in our database, as the recorded targets primarily have smooth surfaces at locations where valid reconstructions are measured. Moreover, our setup minimizes mmWave interactions with external objects, limiting noise primarily to the object itself. As a result, the signal-to-noise ratio of RF ToF is considerably less sensitive to changes in object-to-sensor distance compared to NIR ToF, assuming no interference from external sources.
Figure 9. For selected objects, we show the reconstructed pointclouds (left) next to their deviation from to the MVS reconstruction (right). The signed depth deviation P* is given for each pixel $(u,v)$ in centimeters. All measurements in the domain $\bm{M}^{+}(u,v)$ are projected onto the GT reconstruction and mapped to color using a combination of a symmetrical logarithmic scale and linear mapping between $[-0.5,0.5]$. Furthermore, we quantify the mean deviation of P* in centimeters below each sensor measurement.
7.2. Object-specific Observations
The following section will analyze the objects in Table 4 in regard to their relative depth deviations over one or multiple sensors. For RF ToF, we find that the least deviation relative to the mean of all metrics occurs with planar object geometries (V1 Metal Plate, Cardboard), followed by more complex shapes (Statue, S1 Hand Open, Hand Printed Flat). For a deeper discussion, see Section 9.
Objects made of foam (Sponge), thin plastic (Scrubber, Flowerpot), fabric (Plushie), and paraffin wax (Candle), exhibit the highest depth deviations due to a large fraction of the transmitted RF signal not being immediately reflected. In the case of Mirror, RF penetrates the first (glass) surface and images the silver coating behind, leading to an offset in the depth reconstruction.
For similar reasons, NIR ToF shows large depth deviations for visually transparent objects like Flowerpot, Candle, Sponge, and Tape Dispenser. Although both RF ToF and NIR ToF are susceptible to multi-path effects, our experiments suggest these effects do not occur for the same objects. Further examination of wavelength-specific multi-path effects, with a particular focus on partially transmissive materials, will be discussed in Section 8.
Additional sources of high depth deviation for NIR ToF include thin structures (Scrubber), which reduce the sensor’s effective spatial resolution. Highly reflective objects (Metal Plate) may cause sensor oversaturation, while perfectly specular materials (Mirror) yield depth values not from the object’s surface but from the first weak scatterer encountered along the reflection path.
For the active stereo sensor, we observe higher depth deviations for textureless and partially transmissive materials (Sponge, Candle). Similar to NIR ToF, the uniqueness of the active NIR light pattern can be compromised by multi-path effects. Finally, as expected, the passive stereo camera is particularly sensitive to textureless objects (Sponge, Candle).
8. Discussion ToF in Partially Transmissive Media
As previously discussed in Section 4.2.2, both NIR and RF ToF sensors assume direct reflection and thus are susceptible to internal reflections, such that multi-path effects within the scene may lead to missing or incorrect reconstructions. In this analysis, following the nomenclature by Nayar et al. (2006), we classify radiance transport that involves a single signal bounce between sender and receiver as direct (as, within the sensor’s spatial resolution, it interacts with the scene at one surface point only), and all other types of transport as global (involving multiple scattering or diffraction events within and between objects). Due to their significant difference in wavelength relative to scene features, global radiance transport takes very different forms for each modality. In the case of NIR, representative forms of global transport include inter-reflections, half-transparent surfaces, and subsurface scattering within the object material. Global transport at radio frequencies, on the other hand, is dominated by diffraction and reflections that reshape and redirect the wave front as it interacts with multiple scene elements, and by multiple superimposed responses akin partial transmittance at different depths.
In the remainder, we will now study four selected objects, shown in Figure 9. In addition to Table 4, this figure visualizes depth reconstruction deviations using a signed version P* of our metric P, color-encoded on a symmetrical logarithmic scale (SymLogNorm [^2] https://matplotlib.org/3.8.4/api/_as_gen/matplotlib.colors.SymLogNorm.html), with a linear mapping between $[-0.5,0.5]$ centimeters. The supplementary material includes signed versions of P and $\text{P}_{\text{e}}$ for all MAROON objects.
In the NIR domain, the most prominent effect of global transport occurs for objects with strong internal scattering. Here, the ToF reconstructions exhibit systematic depth deviations of P*, generally biased toward larger distances than the ground truth. This is consistent with the light traveling an additional distance due to scattering within the object before being remitted again, so that the observed propagation time of the actively transmitted signal is consistently longer than for a direct (local) reflection at the object surface. Examples in Figure 9 for internal scattering include subsurface scattering (S1 Hand Open, Sponge, Candle) and inter-reflections within hollow objects (Tape Dispenser). Extended path length due to subsurface scattering is an established effect, systematically measured by Lukinsone et al. (2020). For human skin (as in our dataset S1 Hand Open), and for points of incidence and exitance one millimeter apart, Lukinsone et al. observe effective sub-surface path lengths of up to $26\pm 3$ mm at 800 nm wavelength, which — in the context of a ToF sensor — would result in a systematic depth deviation of half that path length ($\approx+13$ mm). At the same time, however, for human skin a significant portion of the total remitted light stays very close to the point of incidence (Jensen et al., 2001), suggesting that the bulk of the received signal experiences even smaller path length extensions, lending plausibility to our measured systematic depth deviation of $+4.9$ mm for S1 Hand Open to be due to subsurface scattering.
For RF ToF, only one object (Candle) showed a systematic path length extension, suggesting that optical subsurface scattering cannot fully model RF interactions. In contrast to the NIR ToF measurements, the depth deviation for the Candle object is non-uniform, with higher values near the edges due to variations in surface position and orientation that affect radiance transport.
Where the Candle surface faces the antenna array, the received signal is dominated by direct reflections; where direct reflections reflect away from the array (nearer to the candle’s silhouettes), mostly global transport is observed. In accordance with the results by Álvarez López et al. (2018) the depth reconstruction in the parts with little direct reflection appear more distant than ground truth, which the authors attribute to the high relative permittivity $\varepsilon_{r}\approx 2.6$ of paraffin wax that extends the inferred path length under the assumption of speed of light in vacuum.
In summary, objects composed of partially transmissive media primarily yield systematic bias in ToF reconstructions, with estimated depths biased toward larger values than the ground truth. Nevertheless, the factors causing these distortions vary between optical and RF modalities.
9. Discussion of MIMO Radar Signal and Reconstruction
Our observations in Section 7 suggest that RF ToF reconstructions are generally less complete than those of optical sensors, as illustrated in the second row of Figure 9, where quality is best when surface orientations align with the antenna aperture. To identify sources of incomplete RF ToF reconstructions, we first analyze the raw signal response — without inducing additional bias from the reconstruction algorithms — and subsequently relate it to the quality of the measured depth after reconstruction.
In Figure 6, we presented the received signal magnitude across objects with varying material, geometry (surface incidence angle) and size (relative surface area). Following the scatter plot order from left to right, we will now discuss common trends, noting that it remains challenging to disentangle the presented quantities, as the large variability across objects prevents us from isolating one quantity while keeping the others constant. Further results where we discuss the potential relation between object size and object geometry are given in the supplementary material.
Considering object material, metal and metallic-coated objects generally show higher signal magnitudes, which is consistent with previous studies (Ahmed, 2021, 2014). With a considerably lower spread, large magnitudes are also observed for captures of human skin, which is highly reflective due to its rich water content (Ahmed, 2014). Object materials made of polymers, fibers and stone, or foam generally respond with much smaller signal magnitude.
Regarding object geometry and size, no significant global trends are observed; however, consistent patterns emerge within subsets of the same material, particularly in the metal and polymer classes, which have the highest number of samples. For objects of more complex geometry, with a median surface incidence angle greater than $10^{\circ}$, large portions of their surface area face away from the antenna array, resulting in decreased signal responses compared to planar objects aligned closely with the antenna aperture (¡$10^{\circ}$). The reflection direction of the transmitted signal depends on the surface normal’s orientation. As the angle between this normal and the depth axis increases, the solid angle of the object relative to the planar square-shaped antenna aperture (cf. Table 1) decreases. In other words, a decreasing area around the hemisphere of outgoing reflection directions is aligned with the $65^{\circ}$ field of view of the RX antennas, resulting in reduced signal energy reception, and thus radar cross-section (Knott et al., 2004). Superficially this resembles the well-known cosine law in radiometry, but the exact quantitative relationship depends on the object’s location relative to the individual RX and TX antennas and is further modulated by the non-trivial radiation and signal lobes of the antennas.
Aside from object geometry, the received signal magnitude also appears to increase with object size for non-metal materials. Within these material classes, the highest signal magnitude is achieved for objects close to or even larger (objects partially outside FOV) than the antenna aperture. As the latter also typically exhibits a low median surface incidence angle, it remains questionable, whether this observation can be attributed to object size or object geometry. To address this, we additionally visualize the relation between the two quantities in the supplementary material. Assuming that signal magnitude is proportional to the received energy, our findings correspond to the fact that, the reflected signal energy received at the RX antennas directly depends on the surface area of the irradiated object, in case the energy density is constant. To summarize, we identified object geometry and material as two factors that influence the maximum signal magnitude at each receiver, while also observing indications of the relevance of object size.
We now continue to investigate the relationship between signal magnitude and the observed depth deviation, which is illustrated in the upper left scatter plot in Figure 6. Focusing on polymer and metal objects, we generally find no direct relationship between signal magnitude and depth deviation. Polymer and metal objects have a large spread in the $x$- and $y$-axis, respectively, while the opposite axis has comparably low variation. The received signal magnitude may have a more significant influence on the reconstruction quality in less constrained scenarios, involving multiple objects and signal sources, where, for example, depth filtering becomes increasingly relevant. In our experiments, within the RF near field, we suggest that reconstruction quality is more closely related to the distribution of received signals across antennas, influenced primarily by the local antenna layout and characteristics.
Having found no direct relation between signal magnitude and the reconstruction quality, we now examine the depth deviation concerning the previously mentioned three object quantities, which are shown in the bottom row of Figure 6. Similar to our findings for signal magnitude, metal objects generally exhibit the lowest depth deviation with a relatively small spread compared to other material classes, indicating that object material influences reconstruction quality. Furthermore, we find no direct relationship between object size and reconstruction quality. The most notable trend is seen with varying object geometry, where depth error increases alongside the median surface incidence angle across all material classes. The backprojection algorithm assumes that similar energy amounts are received at a point across the majority of RX antennas. Received energy diminishes for surfaces oriented away from the antenna geometry, leading to variations based on antenna positioning. This, in turn, can lead to reduced confidence in the measurements, causing valid data to be filtered out along with noise.
To conclude, the analysis of MIMO imaging radar reconstructions based on the received signal response is considerably less straightforward than that of optical sensors. Our findings suggest that the reconstruction quality of RF ToF sensors is primarily influenced by object geometry, followed by the influence of object material, especially in depth filtering with respect to signal confidence.
9.1. Limitations
Even though the depth deviation of the MIMO imaging radar is on par with that of optical sensors, the reconstructions exhibit significantly more holes, where no valid depth is estimated. While it is intuitive to assume that reconstruction quality is influenced by object material (limiting the returned signal amount — akin to optically transmissive materials), we observe that in our experiments the object geometry is the primary factor of influence and has a greater impact than in optical sensors; however, disentangling the effects of geometry and material remains challenging, as precise impacts on fine-grained surface details cannot be easily assessed. Without a direct mapping between point targets and RX antennas, depth evaluation concerning these surface-level details is infeasible without backprojection. On the other hand, the reconstructed outcomes of backprojection may not align with reality, as the method incorporates a systematic bias by relying on the Born approximation (Ahmed, 2014). To overcome these limitations, we suggest using simulation frameworks, such as employed by Schuessler et al. (2021) or Bräunig et al. (2023a). These could help disentangle potential error sources behind missing reconstructions and enhance the current signal processing methods.
Furthermore, it should be noted that the object reconstructions were evaluated solely for valid locations in the ground-truth data, excluding artifacts like ghost targets or other forms of noise that may arise from violations of the Born approximation. Lastly, we did not capture different orientations of flat objects, which would be an interesting future direction to investigate object orientation in isolation from geometry complexity.
10. Conclusion
In this paper, we present a novel multimodal dataset, MAROON, that allows us to characterize, for the first time, near-field MIMO imaging radars in direct relation with traditional depth imagers from the optical frequency domain for close-range applications. The dataset comprises depth images of a variety of objects, synchronously captured by four mutually calibrated depth imagers and a ground-truth multi-view stereo system. We subsequently analyzed the data within a comprehensive evaluation framework, offering quantitative and qualitative perspectives on each sensor’s depth deviation across multiple metric types, objects, and object-to-sensor distances.
In summary, our analysis revealed the following insights:
  • There is no single sensor modality that would consistently outperform the others. Each sensor has unique strengths and weaknesses related to the object’s material, geometry, and distance from the sensor.
  • NIR ToF displays systematic depth distortions due to effects of global radiance transport within partially transmissive media, whereas RF ToF reconstructions were mostly unaffected.
  • RF ToF partially exhibits missing reconstructions compared to its optical counterpart, primarily due to the object geometry contributing to the reconstruction sparseness.
We hope that the public release of the MAROON dataset will give rise to further study of multimodal sensors in a joint reference frame. The findings presented here are based on aggregate trends and individual object analyses that contribute to the understanding of the addressed sensor characteristics; however, we believe that our dataset still invites further analysis, exploiting the high diversity of the 45 objects that could not be fully addressed in the scope of this paper. Furthermore, we anticipate that insights into the complementary strengths of optical and near-field radio-frequency depth imagers will benefit future sensor fusion applications.